MICHELLE NEEDS YOUR HELP

a) What magnitude point charge creates a 40,000 N/C electric field at a distance of 0.282 m?
1 C

(b) How large is the field at 20.7 m?
2 N/C

Supporting Materials
Answer #1

a)

Electric field due to a charge is given by:

E = k*Q/r^2

where k = 9*10^9

r = distance from the charge Q

So, E = 9*10^9*Q/0.282^2 = 40000

So, Q = 3.53*10^-7 C <------------answer

b)

Now,at r = 20.7 m

E = 9*10^9*(3.53*10^-7)/20.7^2 = 7.41 N/C <-----------answer